# NGO Activism:

### v6

@@ 7 Discussion and Conclusion
Examining NGO campaign timing reveals that not just the decision to intervene, but its timing, can materially affect firm behavior and shape the course of future activism. Over time, NGOs shift from visibility-driven exposure tactics to influence current strategies, highlighting how the temporal dimension of their strategy channels pressure that leads firms to internalize social and environmental externalities.
We first show that NGOs tend to launch their campaigns on the AGM date, likely motivated by the higher media coverage surrounding these events (Result 1). By doing so, NGOs elicit reactions from key stakeholders, such as shareholders. However, the immediate purpose of these campaigns is primarily to gain visibility by voicing stakeholder concerns; they have no direct effect on firms’ current E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
@@
Our results suggest that NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5). First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device. This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following “reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future. [^10] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
Second, stakeholders may come to expect more constructive approaches from established NGOs, such as supporting related shareholder proposals. Consistent with this hypothesis, AGM-date campaigns by mature NGOs—those with multiple prior campaigns—do not generate more negative media coverage than campaigns at other times. By contrast, the very same NGOs do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B10). Hence, we observe not only a shift in strategy as NGOs mature (Figure 4) but also a change in their capacity to inflict reputational harm on firms.
@@
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex) is accounted for through fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied. We leverage the exogeneity of corporate events to uncover a trade-off between visibility and influence that NGOs face when deciding when to campaign, and show that how this trade-off is resolved determines which stakeholders are mobilized (e.g., shareholders today versus shareholders tomorrow). Hence, NGOs act as a coordination device, with distinct implications for the social and environmental impact of targeted firms.
We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to quantitatively study the incentives for collaboration.
@@ Appendix A Omitted Tables
Panel c: Variables varying by target firm and NGO experience
NGO w/ AGM Experience NGO w/o AGM Experience Test difference
@@
Panel d: OLS Regression of $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
@@ Appendix B Omitted Figures
Figure B6: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year. Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in NGO experience over years. Standard errors are clustered at the NGO level.
@@
Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
Figure B8: NGO experience and target firm’s size
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the experience of the NGO, measured by the number of campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 1,338.
Figure B9: NGO experience and target firm’s E&S characteristics
@@
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating the experience of the NGO, measured by either the number of environmental (top panel) or social (bottom panel) campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.
@@ C.1 Staggered Difference-in-Differences Specification
Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1. The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%. By contrast, experienced NGOs see a more modest increase (33%) in search intensity, and only in the month of the campaign.

### v7

@@ 7 Discussion and Conclusion
Examining NGO campaign timing, we show that not just the decision to intervene but also its timing can materially affect firm behavior and shape the course of future activism. Over time, NGOs shift from visibility-driven exposure tactics to those designed to influence corporate strategies, highlighting how the temporal dimension of their actions channels pressure that leads firms to internalize social and environmental externalities.
We first show that NGOs’ early campaigns often coincide with the target’s AGM date, when media attention is highest, and that such timing significantly increases their visibility (Result 1). These visibility gains are concentrated in NGOs’ first few campaigns, indicating that AGM-day actions serve primarily to build name recognition rather than to influence firms directly. By contrast, later AGM-day campaigns have no immediate effect on firms’ E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
@@
Our results suggest that NGOs face a trade-off that they navigate by adapting their strategies over time. NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
As a result, having targeted AGMs in the past increases the likelihood that an NGO can help a shareholder proposal pass related to its campaign. However, repeatedly targeting AGMs leads to weaker influence over firm behavior (Result 5).
One interpretation is that NGOs, having already demonstrated commitment to their cause, face weaker incentives to keep signaling through costly AGM-day campaigns. In this view, declining returns simply reflect a reduced need for visibility once reputational capital is established.
Yet the evidence points to a different mechanism. Within the same NGO, AGM-day campaigns progressively lose their marginal effect on media attention (Figure 2) and their ability to damage a firm’s reputation (Appendix Figure B10), suggesting that the audience’s response—rather than the NGO’s effort—changes over time. As NGOs become recognized and credible actors, stakeholders interpret their campaigns differently: what initially signals toughness later signals redundancy or routine. Established NGOs are expected to engage in more constructive dialogue, such as supporting shareholder proposals, rather than resorting to public confrontation. Consequently, the same action produces less reputational harm for firms and fewer visibility gains for NGOs.
The diminishing returns to AGM targeting thus stem from evolving stakeholder expectations and a changing perception of NGO legitimacy, rather than from reduced internal pressure to seek exposure. [^11] This interpretation does not rule out signaling motives altogether. Early AGM-day campaigns are consistent with “reputation-for-toughness” models, in which actors bear short-term costs to establish credibility that can be leveraged in future interactions (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
@@
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex) or experience in managing AGM-related information is accounted for through NGO-by-year fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
In conclusion, NGOs without an established reputation coordinate stakeholders over time by campaigning on AGM dates, drawing attention to issues that spur future engagement. Once reputable, they coordinate stakeholders within the same AGM cycle by supporting existing shareholder proposals and directly shaping firm decisions. Thus, NGOs act as a coordination device throughout their lifecycle, with the form of coordination evolving with their perceived toughness and credibility. In this way, NGOs’ incentives shape their strategies and, ultimately, the social and environmental impact of firms.
We also acknowledge that NGOs increasingly engage not only in campaigns but also in collaborations with corporations (Dahan et al., 2010). Future research could examine when confrontation shifts to cooperation and how such collaboration reshapes the balance between visibility, credibility, and influence.
@@ Appendix A Omitted Tables
Panel c: Selection—Avg. target characteristics’ based on whether the campaigning NGO campaigned on AGM dates in the past
NGO w/ Past AGM Campaigns NGO w/o Past AGM Campaigns Test Difference
@@
Panel d: Selection—Target’s characteristics based on NGO’s past campaigns regressing $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{\# Past Campaign}_{g(n)y}+\varepsilon_{ny}$
@@
  • Note: This table reports summary statistics and robustness checks across four panels.
  • Panel (a). Summary statistics on NGO campaigns (Sigwatch); AGM dates (ISS).
  • Panel (b). Target-firm statistics: accounting (Compustat), ESG (MSCI), news counts (RavenPack).
  • Panel (c). No-selection-on-observables test comparing targets where the NGO had previously campaigned on at least one of its targets’ AGM dates. Market Cap, Total Assets, and Sales are deciles using all U.S. Compustat firms in the campaign quarter. Deciles avoid mechanical correlation with # Past NGO Campaigns, which grows over time.
@@ Appendix B Omitted Figures
Figure B6: NGO past campaigns and the timing of current NGO campaigns – Alternative set of Fixed Effects
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the number of campaigns the focal NGO had already launched. The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year. Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in the numbers of NGO campaigns over years. Standard errors are clustered at the NGO level.
@@
Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the number of past campaigns of the focal NGO. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
Figure B8: Past campaigns and target firm’s size
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the number of campaigns the focal NGO had launched in previous years as in (5). Standard errors are clustered at the firm level. The number of observations is 1,338.
Figure B9: Past campaigns and target firm’s E&S characteristics
@@
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating either the number of environmental (top panel) or social (bottom panel) campaigns the focal NGO had launched in previous years as in (5). Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.
@@ C.1 Staggered Difference-in-Differences Specification
Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1. The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%. By contrast, NGOs see a more modest increase (33%) in search intensity for later campaigns, and only in the month of the campaign.